# Color-superconducting quarkyonic matter

### v1

@@ IV.2 Conformal limit
As shown in Fig. 5, CSQY matter asymptotically approaches the conformal limit of QCD, characterized by $c_{S}^{2}=1/3$ and $\delta=0$ Kurkela et al. (2010); Fraga et al. (2014); Gorda et al. (2018); Fernandez and Kneur (2022). Consistent with perturbative QCD results, these conformal values of the speed of sound and the dimensionless interaction measure are approached from below and above, respectively. To clarify this feature of the present model, we would like to consider its high-density asymptote.
@@
For the sake of simplicity, we consider the symmetric case, which preserves the qualitative conclusions of the analysis. In this regime, the baryonic chemical potential determines the thermodynamic properties of CSQY matter. A small deviation from the conformal behaviour is caused by the presence of three-dimensional scales, i.e., the parameter $\Lambda_{a}$ controlling the thickness of the nucleon shell and in the symmetric case coinciding with $\Lambda$, pairing gap $\Delta$, and quark masses $m_{q}$. To account for the first of them, we notice that the vanishing of the nucleon shell thickens at high chemical potentials brings their number densities to the form $n_{a}=\Lambda^{\kappa+1}k_{a}^{2-\kappa}/\pi^{2}$, where the maximum momentum of nucleons $k_{a}$ should be replaced by its asymptote $\mu_{a}(\kappa-2)/(\kappa-3)$ found as the leading order expansion of Eq. (III.1). The effects of quark pairing can be perturbatively accounted for by using the corresponding correction to the quark pressure of one color-flavor state $\Delta_{fc}^{2}\mu_{fc}^{2}/4\pi^{2}$, where $\Delta_{fc}=(\Delta_{\mu},\Delta_{\mu},0)$ and $\Delta_{\mu}$ is the momentum-dependent pairing gap evaluated at $|{\bf k}|=\mu_{B}/3$ (see Section II.C of Ref. Blaschke et al. (2023b) for a derivation). It is worth mentioning that this correction coincides with the pairing gap term of the well-known phenomenological parameterization of the quark matter EoS by Alford, Braby, Paris, and Reddy Alford et al. (2005) if the gap $\Delta_{\mu}$ is treated as a medium-independent constant. Finally, the mass correction to the quark number density can be obtained by using the leading order expansion of the number density of unpaired quarks with the quark Fermi momenta expressed through their chemical potential as $k_{fc}=\mu_{fc}-m_{q}^{2}/2\mu_{fc}$. Thus, the high-density asymptote of the quark number density becomes $n_{fc}=\mu_{fc}^{3}/3\pi^{2}-\mu_{fc}m_{q}^{2}/2\pi^{2}-\mu_{fc}^{3}\Delta_{fc}^{2}/\Lambda^{2}\pi^{2}$, where the quark pairing term was obtained by differentiating the corresponding pressure correction with respect to $\mu_{fc}$ and keeping the leading order contribution only. To account for the condition of color neutrality (29), we express the quark chemical potentials as $\mu_{fc}=\mu_{B}/3+\mu_{8}\lambda_{8c}$ and notice that $\mu_{8}$ is small, since $\Delta_{fc}\rightarrow 0$ at high densities. Consequently, linearizing Eq. (29) in $\mu_{8}$, we find the later as $\mu_{8}=\mu_{B}\Delta_{\mu}^{2}/9\Lambda^{2}$. It is important to stress that the above decomposition of the quark chemical potential, along with the conditions of strong equilibrium, leads to $\mu_{a}=\mu_{B}$. Using the above results, the leading order expression for the baryon density of symmetric CSQY matter at high densities becomes
@@
With this expression, the asymptotic behaviour of the speed of sound can be found as
@@
The asymptote of the dimensionless interaction measure $\delta=1/3-c_{S}^{2}$ is obtained using the L’Hôpital’s rule Ivanytskyi (2025a, b).
@@ IV.3 Chemical composition
Fig. 6 shows the fractions of the considered particle species as a function of baryon density. The qualitative behavior is similar to the results of Refs. Zhao and Lattimer (2020); Margueron et al. (2021). Before the quark onset is reached, only protons, neutrons, and electrons are present. As the baryon density increases, the fraction of protons and electrons rises until the quark onset density is reached. At $n_{B}^{\rm onset}$, nucleons begin to dissolve into quarks, resulting in a substantial increase of the quark fractions. We do not discriminate between paired and unpaired quarks, only showing the species of each flavor. Clearly, the number of paired (green and red) quarks is twice the number of unpaired (blue) ones. The two panels of Fig. 6 show the particle fractions for two different values of the quark onset density.
@@
The particle abundances strongly affect a star’s thermal evolution and neutrino emission Page et al. (2006); Ávila et al. (2024). Among the various processes occurring in the NS interior, the direct Urca (DU) process, corresponding to direct and inverse $\beta$-decays of neutrons in nuclear matter, is considered as the most efficient cooling mechanism. The onset of this process is determined by the proton fraction, which, once sufficiently large, leads to enhanced neutrino emission and rapid stellar cooling. In the present model, the maximum proton fraction is determined solely by the hadronic EoS. However, as shown by Grigorian et al. (2005); Sagun et al. (2023), quark DU processes can significantly affect the cooling behavior and highlight the role of CS in NS thermal evolution. We leave the detailed analysis of this aspect for the following study.
@@
Figure 8: The tidal deformability as a function of the gravitational mass corresponding to the same EoSs as in Fig. 4. The orange region depicts the constraint obtained from GW170817 Abbott et al. (2018). The purely hadronic configurations without an onset of QY matter are illustrated by the black dash-dotted curve in both panels.
@@ IV.4 Color-superconducting quarkyonic stars
Solving the Tolman–Oppenheimer–Volkoff equations Tolman (1939); Oppenheimer and Volkoff (1939), we obtain static, spherically symmetric equilibrium configurations of NSs with CSQY cores. Fig. 7 presents the mass–radius relation (see the left column) and the mass–central baryon density relation (see the right column). The purely hadronic sequences are shown by the black dash-dotted curves. The onset of configurations containing CSQY cores is indicated by the open black circles, with colored curves that emerge from this point, indicating the CSQY branches. The open red circle denotes the onset of causality violation in the purely hadronic case. Each sequence is plotted up to the maximum mass configurations. It is seen that the onset of CSQY matter leads to higher radii and masses of NSs, resulting in a more pronounced back-bending of the mass-radius relations.
@@
As discussed earlier, larger values of $\kappa$ and the effective pairing gap $\Delta$ systematically result in higher NS masses and radii. The strength of these effects depends on the onset density and on the stiffness of the hadronic EoS (see the left column of Fig. 7). The onset of CSQY matter at high densities diminishes the impact of quark degrees of freedom on NS properties, owing to the comparatively small size of the resulting CSQY cores. A similar effect arises from the stiffness of the hadronic EoS. As a result, for later onsets, the role of the hadronic EoS becomes increasingly dominant, allowing its properties to be more tightly constrained within the model. In general, our main results show a good agreement with the existing astrophysical and gravitational wave constraints. The transition to CSQY matter allows for more massive NSs, while also predicting significantly larger radii for high-mass stars containing a CSQY core. This is an important result, as pulsars heavier than $2.2\leavevmode\nobreak\ M_{\odot}$ are conceivable Romani et al. (2022).
@@
Fig. 8 shows the behavior of the tidal deformability $\Lambda$ as a function of the NS mass. In the case of the early onset of CSQY matter at $n_{B}^{\text{onset}}=2n_{0}$ the larger NS radii discussed earlier lead to higher values of $\Lambda$ compared to the purely hadronic case (shown by the black dash-dotted curve), since $\Lambda\propto(R/M)^{5}$ Hinderer (2008); Damour and Nagar (2009). For the later onset at $n_{B}^{\text{onset}}=3n_{0}$e have explored a separate, dynamical signature of first-order phase transitions: the resonant tidal excitation of interfacial i -modes. The detection of an interface mode in a coalescing neutron-star binary would be a smoking-gun signature of a first-order phase transition, distinguishable in a single gravitational-wave event. , the deviation from the purely hadronic curve is small, with only a slight decrease in $\Lambda$ at high masses. The results are compared with the 90% confidence-level region inferred from the gravitational-wave event GW170817 Abbott et al. (2018).
Refer to captionRefer to caption
Figure 9: The same as Fig. 4, but obtained with the soft nuclear EoS.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ V Conclusions
The developed model of NS matter with CSQY cores provides a good agreement with the observational data. At the same time, a more systematic analysis of its parameter space is required. The present model does not account for the effects of spontaneous breaking and dynamical restoration of chiral symmetry, and the effects of vector repulsion at high densities. Studying them along with using a self-consistently defined pairing gap deserves a separate study.
@@ Acknowledgments
The authors thank S. Reddy, A. Schmitt, and D. Blaschke for the fruitful discussions and valuable comments. C.G. and I.L. express their gratitude to the Fundação para a Ciência e Tecnologia (FCT), Portugal, for providing financial support to the Center for Astrophysics and Gravitation (CENTRA/IST/ULisboa) through Grant Project No. UIDB/00099/2025. C.G. also acknowledges the Fundação para a Ciência e Tecnologia (FCT), Portugal, through the IDPASC PT-CERN program with the No. PRT/BD/154664/2022. This work of O.I. was performed within the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science and received funding from the Polish National Science Center under grant No. 2021/43/P/ST2/03319. V.S. gratefully acknowledges support from the UKRI-funded “The next-generation gravitational-wave observatory network” project (Grant No. ST/Y004248/1). This work was produced with the support of INCD and funded by FCT I.P. under Advanced Computing Project 2023.10526.CPCA.A2 with DOI identifier 10.54499/2023.10526.CPCA.A2.
@@
The data supporting the findings of this article are openly available, provided that an appropriate citation is given Gärtlein et al (2025).
@@ Appendix A Soft hadronic EoS
As discussed in Sec. III, we address the role of the nuclear mater EoS by considering here its softer version. The latter is obtained within the same metamodeling approach as the hadronic EoS in Sec. III, but suppressing the term corresponding to the skewness in the isoscalar interaction potential, i.e., $Q_{0}=0$ in Eq. (III.1), while the rest of the nuclear matter parameters remain equal to the ones in Table. 1. As a result, the hadronic part of the CSQY EoS significantly softens. In Fig. 9, the pressure as a function of energy density shows, as expected, the black dash-dotted curve to be below the heavy pulsars constraints.
@@
The speed of sound squared depicted in Fig. 10 shows a pronounced peak after the onset of quarks. Interestingly, the maximum values of $c_{s}^{2}$ are increased compared to the case of a stiff hadronic EoS. Thus, the peak at the onset violates causality for $\kappa=9$ and $\Delta\neq 0$, with values of $c_{s}^{2}>1$. Therefore, the combinations of $\kappa=9$ and $\Delta>0$ in the case of the soft hadronic EoS are excluded from the realistic parameters of the CSQY EoSs. Conformality is still reached from below in all cases, as seen in Fig. 10. In comparison to Fig. 5, consistency with the illustrated constraints Altiparmak et al. (2022); Marczenko et al. (2023) is provided for the early onset case.

### v2

@@ IV.2 Conformal limit
As shown in Fig. 5, CSQY matter asymptotically approaches the conformal limit of QCD, characterized by $c_{S}^{2}=1/3$ and $\delta=0$ Kurkela:2009gj; Fraga:2013qra; Gorda:2018gpy; Fernandez:2021jfr. Consistent with perturbative QCD results, these conformal values of the speed of sound and the dimensionless interaction measure are approached from below and above, respectively.
As discussed in Ref. Blaschke:2022egm, nonvanishing current masses of quarks and repulsive interaction among them is the physical reason of reaching the conformal limit as described above. The high density asymptotes of $c_{S}^{2}$ and $\delta$ found within the three-flavor nonlocal NJL model confirm this conclusion Ivanytskyi:2024zip. To clarify the features of the speed of sound and the dimensionless interaction measure in the present model, we analyze its high-density asymptotic behavior.
@@
For the sake of simplicity, we consider the symmetric case, which preserves the qualitative conclusions of the analysis. In this regime, the baryonic chemical potential determines the thermodynamic properties of CSQY matter. A small deviation from the conformal behavior is caused by the presence of three-dimensional scales, i.e., the parameter $\Lambda_{a}$ controlling the thickness of the nucleon shell and in the symmetric case coinciding with $\Lambda$, pairing gap $\Delta$, and quark masses $m_{q}$. To account for the first of them, we notice that the vanishing of the nucleon shell thickens at high chemical potentials brings their number densities to the form $n_{a}=\Lambda^{\kappa+1}k_{a}^{2-\kappa}/\pi^{2}$, where the maximum momentum of nucleons $k_{a}$ should be replaced by its asymptote $\mu_{a}(\kappa-2)/(\kappa-3)$ found as the leading order expansion of Eq. (III.1). The effects of quark pairing can be perturbatively accounted for by using the corresponding correction to the quark pressure of one color-flavor state $\Delta_{fc}^{2}\mu_{fc}^{2}/4\pi^{2}$, where $\Delta_{fc}=(\Delta_{\mu},\Delta_{\mu},0)$ and $\Delta_{\mu}$ is the momentum-dependent pairing gap evaluated at $|{\bf k}|=\mu_{B}/3$ (see Section II.C of Ref. Blaschke:2022egm for a derivation). It is worth mentioning that this correction coincides with the pairing gap term of the well-known phenomenological parameterization of the quark matter EoS by Alford, Braby, Paris, and Reddy Alford:2004pf if the gap $\Delta_{\mu}$ is treated as a medium-independent constant. Finally, the mass correction to the quark number density can be obtained by using the leading order expansion of the number density of unpaired quarks with the quark Fermi momenta expressed through their chemical potential as $k_{fc}=\mu_{fc}-m_{q}^{2}/2\mu_{fc}$. Thus, the high-density asymptote of the quark number density becomes $n_{fc}=\mu_{fc}^{3}/3\pi^{2}-\mu_{fc}m_{q}^{2}/2\pi^{2}-\mu_{fc}^{3}\Delta_{fc}^{2}/\Lambda^{2}\pi^{2}$, where the quark pairing term was obtained by differentiating the corresponding pressure correction with respect to $\mu_{fc}$ and keeping the leading order contribution only. To account for the condition of color neutrality (29), we express the quark chemical potentials as $\mu_{fc}=\mu_{B}/3+\mu_{8}\lambda_{8c}$ and notice that $\mu_{8}$ is small, since $\Delta_{fc}\rightarrow 0$ at high densities. Consequently, linearizing Eq. (29) in $\mu_{8}$, we find the later as $\mu_{8}=\mu_{B}\Delta_{\mu}^{2}/9\Lambda^{2}$. It is important to stress that the above decomposition of the quark chemical potential, along with the conditions of strong equilibrium, leads to $\mu_{a}=\mu_{B}$. Using the above results, the leading order expression for the baryon density of symmetric CSQY matter at high densities becomes
@@
With this expression, the asymptotic behavior of the speed of sound can be found as
@@
The asymptote of the dimensionless interaction measure $\delta=1/3-c_{S}^{2}$ is obtained using the L’Hôpital’s rule Ivanytskyi:2024zip; Ivanytskyi:2025cnn.
@@ IV.3 Chemical composition
Fig. 6 shows the fractions of the considered particle species as a function of baryon density. The qualitative behavior is similar to the results of Refs. Zhao:2020dvu; Margueron:2021dtx. Before the quark onset is reached, only protons, neutrons, and electrons are present. As the baryon density increases, the fraction of protons and electrons rises until the quark onset density is reached. At $n_{B}^{\rm onset}$, nucleons begin to dissolve into quarks, resulting in a substantial increase of the quark fractions. We do not discriminate between paired and unpaired quarks, only showing the species of each flavor. Clearly, the number of paired (green and red) quarks is twice the number of unpaired (blue) ones. The two panels of Fig. 6 show the particle fractions for two different values of the quark onset density.
@@
The particle abundances strongly affect a star’s thermal evolution and neutrino emission Page:2005fq; Avila:2023rzj. Among the various processes occurring in the NS interior, the direct Urca (DU) process, corresponding to direct and inverse $\beta$-decays of neutrons in nuclear matter, is considered as the most efficient cooling mechanism. The onset of this process is determined by the proton fraction, which, once sufficiently large, leads to enhanced neutrino emission and rapid stellar cooling. In the present model, the maximum proton fraction is determined solely by the hadronic EoS. However, as shown by Grigorian:2004jq; Sagun:2023rzp, quark DU processes can significantly affect the cooling behavior and highlight the role of CS in NS thermal evolution. We leave the detailed analysis of this aspect for the following study.
@@
Figure 8: The tidal deformability as a function of the gravitational mass corresponding to the same EoSs as in Fig. 4. The orange region depicts the constraint obtained from GW170817 LIGOScientific:2018cki. The purely hadronic configurations without an onset of QY matter are illustrated by the black dash-dotted curve in both panels.
@@ IV.4 Color-superconducting quarkyonic stars
Solving the Tolman–Oppenheimer–Volkoff equations Tolman:1939jz; Oppenheimer:1939ne, we obtain static, spherically symmetric equilibrium configurations of NSs with CSQY cores. Fig. 7 presents the mass–radius relation (see the left column) and the mass–central baryon density relation (see the right column). The purely hadronic sequences are shown by the black dash-dotted curves. The onset of configurations containing CSQY cores is indicated by the open black circles, with colored curves that emerge from this point, indicating the CSQY branches. The open red circle denotes the onset of causality violation in the purely hadronic case. Each sequence is plotted up to the maximum mass configurations. It is seen that the onset of CSQY matter leads to higher radii and masses of NSs, resulting in a more pronounced back-bending of the mass-radius relations.
@@
As discussed earlier, larger values of $\kappa$ and the effective pairing gap $\Delta$ systematically result in higher NS masses and radii. The strength of these effects depends on the onset density and on the stiffness of the hadronic EoS (see the left column of Fig. 7). The onset of CSQY matter at high densities diminishes the impact of quark degrees of freedom on NS properties, owing to the comparatively small size of the resulting CSQY cores. A similar effect arises from the stiffness of the hadronic EoS. As a result, for later onsets, the role of the hadronic EoS becomes increasingly dominant, allowing its properties to be more tightly constrained within the model. In general, our main results show a good agreement with the existing astrophysical and gravitational wave constraints. The transition to CSQY matter allows for more massive NSs, while also predicting significantly larger radii for high-mass stars containing a CSQY core. This is an important result, as pulsars heavier than $2.2\penalty 10000\ M_{\odot}$ are conceivable Romani:2022jhd.
In general, the proposed CSQY EoS provides a good description of the observational data. The hadronic curves in Fig. 7 also agree well with the NS constraints. However, the used hadronic EoS does not include hyperons. Accounting for them would significantly soften the EoS and make it barely reach the two solar mass configurations (see e.g. Refs. Yamamoto:2015lwa; Yamamoto:2017wre; Li:2019sxd for discussion of this effect), which in turn would lead to a loss of consistency with the PSR J0740+6620 constraint Salmi:2024aum. As it was discussed in Refs. Shahrbaf:2019vtf; Yamamoto:2021htv, transition to stiff quark matter, which in this work is realized within the CSQY framework, resolves this tension of the purely hadronic scenario. All the depicted mass-radius curves agree with the PSR J0614-3329 data Mauviard:2025dmd at the 92% confidence level. The same quality of the data description is reached with respect to the HESS J1731-347 object Doroshenko2022. Further improvement of the agreement with the low mass constraints from PSR J0614-3329 and HESS J1731-347 requires a more elaborate hadronic EoS and a Bayesian scan over the parameters of the CSQY matter EoS, which is beyond the scope of the present work.
@@
Fig. 8 shows the behavior of the tidal deformability $\Lambda$ as a function of the NS mass. In the case of the early onset of CSQY matter at $n_{B}^{\text{onset}}=2n_{0}$ the larger NS radii discussed earlier lead to higher values of $\Lambda$ compared to the purely hadronic case (shown by the black dash-dotted curve), since $\Lambda\propto(R/M)^{5}$ Hinderer:2007mb; Damour:2009vw. For the later onset at $n_{B}^{\text{onset}}=3n_{0}$e have explored a separate, dynamical signature of first-order phase transitions: the resonant tidal excitation of interfacial i -modes. The detection of an interface mode in a coalescing neutron-star binary would be a smoking-gun signature of a first-order phase transition, distinguishable in a single gravitational-wave event. , the deviation from the purely hadronic curve is small, with only a slight decrease in $\Lambda$ at high masses. The results are compared with the 90% confidence-level region inferred from the gravitational-wave event GW170817 LIGOScientific:2018cki.
@@ V Conclusions
The developed model of NS matter with CSQY cores provides a good agreement with the observational data. At the same time, a more systematic analysis of its parameter space is required. The model does not account for the effects of spontaneous breaking and dynamical restoration of chiral symmetry, and the effects of vector repulsion at high densities. Studying them along with using a self-consistently defined pairing gap deserves a separate study.
@@ Acknowledgments
The authors thank S. Reddy, A. Schmitt, and D. Blaschke for the fruitful discussions and valuable comments. C.G. and I.L. express their gratitude to the Fundação para a Ciência e Tecnologia (FCT), Portugal, for providing financial support to the Center for Astrophysics and Gravitation (CENTRA/IST/ULisboa) through Grant Project No. UIDB/00099/2025. C.G. also acknowledges the Fundação para a Ciência e Tecnologia (FCT), Portugal, through the IDPASC PT-CERN program with the No. PRT/BD/154664/2022. This work of O.I. was performed within the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science and received funding from the Polish National Science Center under Grant No. 2021/43/P/ST2/03319. V.S. gratefully acknowledges support from the UKRI-funded “The next-generation gravitational-wave observatory network” project (Grant No. ST/Y004248/1). This work was produced with the support of INCD and funded by FCT I.P. under Advanced Computing Project No. 2023.10526.CPCA.A2 with DOI identifier 10.54499/2023.10526.CPCA.A2.
[…]
Figure 9: The same as Fig. 4, but obtained with the soft nuclear EoS.
@@
The data supporting the findings of this article are openly available, provided that an appropriate citation is given gärtlein_2025_170380983.
Refer to captionRefer to caption
@@ Appendix A Soft hadronic EoS
As discussed in Sec. III, we address the role of the nuclear mater EoS by considering here its softer version. The latter is obtained within the same metamodeling approach as the hadronic EoS in Sec. III, but suppressing the term corresponding to the skewness in the isoscalar interaction potential, i.e., $Q_{0}=0$ in Eq. (III.1), while the rest of the nuclear matter parameters remain equal to the ones in Table. 1. As a result, the hadronic part of the CSQY EoS significantly softens. In Fig. 9, the pressure as a function of energy density shows, as expected, the black dash-dotted curve to be below the heavy pulsar constraints.
@@
The speed of sound squared depicted in Fig. 10 shows a pronounced peak after the onset of quarks. Interestingly, the maximum values of $c_{s}^{2}$ are increased compared to the case of a stiff hadronic EoS. Thus, the peak at the onset violates causality for $\kappa=9$ and $\Delta\neq 0$, with values of $c_{s}^{2}>1$. Therefore, the combinations of $\kappa=9$ and $\Delta>0$ in the case of the soft hadronic EoS are excluded from the realistic parameters of the CSQY EoSs. Conformality is still reached from below in all cases, as seen in Fig. 10. In comparison to Fig. 5, consistency with the illustrated constraints Altiparmak:2022bke; Marczenko:2022jhl is provided for the early onset case.